OpenStudy (anonymous):

A chemist needs to mix a 12% acid solution with a 32% acid solution to obtain an 8-liters mixture consisting of 20% acid. How many liters of each of the acid solutions must be used?

OpenStudy (anonymous):

x + y = 8 0.12x + 0.32 y = 8(0.20)

OpenStudy (anonymous):

Can you solve the system?

OpenStudy (anonymous):

can you tell me what I need to do first?

OpenStudy (anonymous):

y = 8-x 0.12x+ 0.32(8-x) = 1.6

OpenStudy (anonymous):

I solved the simple equation to y = 8-x and then substituted into the 2nd equation

OpenStudy (anonymous):

After you simplify, that will tell you how much of the 12% solution you need. You'll have to then solve for the 32% solution by subtracting from 8

OpenStudy (anonymous):

so, 4.8 for the 12%?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then what do I do?

OpenStudy (anonymous):

u need the 32% solution amount so 8 - 4.8 = ???

OpenStudy (anonymous):

3.2

OpenStudy (anonymous):

yes...be sure to label your answers